OpenStudy (rayford):

Identify the horizontal asymptote of f(x)= 4x/7

OpenStudy (cwrw238):

are you sure you have question right?

OpenStudy (rayford):

one moment i'll draw it :)

OpenStudy (rayford):



OpenStudy (cwrw238):

this equation represents a straight line through the origin

OpenStudy (rayford):

well here are the choices..

OpenStudy (rayford):

y=4/7 y=0 y=7/4 no horizontal asytope

OpenStudy (rayford):

soo...it's no horizontal?

OpenStudy (cwrw238):

yes

OpenStudy (rayford):

just making sure thank you very much.

OpenStudy (cwrw238):

yw

OpenStudy (rayford):

and thank you cuzzin for last time :)

OpenStudy (anonymous):

Yes, no horizontal. For a horizontal asymptote to occur, either (1) the degree of the numerator is less than that of the denominator (in which case, HA is y=0), or (2) the degrees in the numerator and denominator are equal, in which case you would divinde the leading coefficient in the numerator by the leading coefficient in the denominator. This will give you your HA.

OpenStudy (anonymous):

No prob

OpenStudy (rayford):

you guys are life savers.. thank you so much